# Dub-S2ST: Textless Speech-to-Speech Translation for Seamless Dubbing

### v1

@@ 4.1 Dataset
CVSS-C Jia et al. (2022b) is a widely-used dataset for speech-to-speech translation, consisting of 21 languages to English translations where English speech is generated by single-speaker TTS model. The proposed method is trained and evaluated with French-English (fr-en) subset, due to the abundance of samples compared to other language pairs. French-English subset contains 207,364 (train) / 14,759 (dev) / 14,759 (test) pairs of source and target speech samples, totaling 264 hours. The experiment utilizes train and dev split for training and validation, and test split for final evaluation.
@@ 4.2 Implementation Details
Preprocessing. Audio samples are resampled to 16kHz and preprocessed with Voice Activity Detection (VAD) toolhttps://github.com/snakers4/silero-vad to remove unnecessary silence and paralinguistic information at the beginning and end. The samples are then processed with pretrained mHuBERT Lee et al. (2021) and k-means clustering model to obtain discrete unitshttps://github.com/facebookresearch/fairseq/blob/main/examples/speech_to_speech/docs/textless_s2st_real_data.md.
Architecture. To maintain consistency with prior works on speech-to-unit translation models, we design our model with 12 Conformer encoder layers and 6 DiT layers. The total model takes 61M parameters. The unit-to-speech synthesizer consists of 4 Down, Mid, and UpBlock, where each is a cascade of 1D Convolution and Transformer layer.
Training. The speech-to-unit translation model was trained with a total batch size of 3,200 seconds for 100k updates. We used 8 RTX A5000 GPUs for training, and the total training took approximately 10 hours. We optimized the models with the AdamW Loshchilov and Hutter (2019) optimizer and applied a dropout rate of 0.3. The learning rate was warmed up for the first 10k steps to a peak of $1\times 10^{-3}$, and then decayed using an inverse square root schedule. We implemented our approach using Fairseq Ott et al. (2019). The unit-to-speech module was initialized from Du et al. (2024) and fine-tuned with fixed learning rate of $1\times 10^{-4}$ for 200k steps using LRS3 Afouras et al. (2018) dataset, an English multi-speaker corpus.
@@ 4.3 Evaluation Metrics
ASR-BLEU Lee et al. (2022) is a widely adopted metric for assessing S2ST quality. It measures translation quality by transcribing the generated speech using a pretrained ASR model Baevski et al. (2020) and comparing it with the ground-truth text to compute BLEU Post (2018) https://github.com/facebookresearch/fairseq/tree/ust/examples/speech_to_speech/asr_bleu.
@@
S2UT Lee et al. (2022) 24.54 3.784 0.036 3.922
CTC-S2UT Fang et al. (2024) 24.51 3.785 0.037 3.908
UTUT Kim et al. (2024) 26.49 3.840 0.036 3.927
w/ Zero-shot Vocoder Choi et al. (2024) 26.33 3.882 0.145 3.101
[…]
DiffNorm Tan et al. (2024a) 19.53 - - -
@@
BLASER 2.0 Dale and Costa-jussà (2024) serves as an automatic measure for assessing semantic similarity between source and generated speech. We adopt its reference-free variant, BLASER 2.0-QEhttps://huggingface.co/facebook/blaser-2.0-qe, which estimates translation quality without reference text unlike BLEU score.
SIM evaluates the speaker similarity between the generated and original speech. We use a pretrained speaker verification model, a fine-tuned version of WavLM-Large Chen et al. (2022), to extract speaker embedding vectors and calculated cosine similarity between the two.
DNSMOS (Deep Noise Suppression Mean Opinion Score) Reddy et al. (2021) is an automated perceptual speech quality assessment of generated speechWe used a model trained with ITU-T P.808: https://github.com/microsoft/DNS-Challenge/tree/master/DNSMOS. The quality is estimated with a score in the range of [1, 5] where the larger value indicates higher speech quality.
Duration Compliance (DC) Wu et al. (2023) calculates the portion of generated speech whose duration ratio with the source lies within certain range. This indicates how S2ST system preserves the duration when generating translated speech.
Speed Compliance (SC) represents a similar trend to LC, but operates on the ratio between the speed of speech, measured by syllable per secondhttps://github.com/facebookresearch/seamless_communication/blob/main/docs/expressive/README.md. This measure shows how much the speed of generated speech lies in that of the source.
@@ 5.1 Quantitative Comparison
Duration and Speed Analysis. We evaluate our model’s performance in duration control by measuring the generated speech duration relative to the source, as shown in Table 2. In addition to standard S2ST baselines, we compare against methods that incorporate duration control: VideoDubber Wu et al. (2023), which employs additional positional embeddings, and TransVIP Le et al. (2024), which introduces isochrony positional embedding based on voice activity information. While existing methods achieve reasonable compliance within a 40% threshold, they struggle under a stricter 20% constraint. In contrast, our model achieves exact duration matching through explicit length initialization in the decoder, yielding 100% compliance.
The proposed model’s advantage becomes more evident in the speed analysis, where our model, both with and without speed adaptation, outperforms all baselines in speed compliance. This indicates that, when conditioned on the source duration, our model not only matches the duration but the speaking pace of the source. The proposed speed adaptation further enhances this alignment. Overall, these results underscore the robustness of our approach in controlling speech duration and speed, both of which serve as critical factors in dubbing.
@@ 5.2 Human Evaluation
The speed consistency evaluation confirms that our proposed speed adaptation method effectively aligns the speaking speed between source and target speeches, producing translated speech at a pace closely matching the source. This implicit synchronization contributes to improve the translation consistency, since it helps the model generate sequences in similar speed to the source when given its duration, minimizing unnecessary pauses and repetitive words. Notably, significant differences in speech naturalness emerged during the evaluation. Baseline samples required manual waveform adjustments, leading to considerable degradation in perceptual quality. Conversely, our proposed approach yielded accurate, time-aligned speech outputs, eliminating the need for post-processing adjustments and maintaining high naturalness suitable for dubbing applications.
@@ 5.3 Discussion
Can explicit duration control truly change the translated content? As illustrated in Figure 3, our model responds to different duration prompts by producing semantically consistent outputs with nuanced variations in structure and phrasing. For example, “none of this happened” is rendered as “nothing took place” at 80% duration and “nothing of all that took place” at 120%, each expressing the same meaning without repetition or loss of information. Moreover, the relative number of characters of the translated speeches with different durations in Table 5 validates that the model flexibly generates semantics that fit the given duration, rather than forcing a fixed translation to fit varying lengths.
@@ 5.4 Ablation Study
Model architecture and Masking schedule. To assess how different model architectures and masking schedules affect the speech-to-unit translation model of Dub-S2ST, we compare several design choices, as shown in Table 6. The last row presents ASR-BLEU performances from Dub-S2ST-single without speed adaptation, evaluated across varying numbers of function evaluations (NFE). Replacing our DiT decoder with a standard Transformer decoder consistently degrades translation quality across all NFEs, indicating that incorporating diffusion timestep into the model benefits translation learning. We also examine the effect of the masking schedule by comparing our linear schedule with a cosine schedule. The results show that the linear schedule yields overall better performance, suggesting its effectiveness without losing translation quality. Based on the latency-performance trade-off, we opt to choose the NFE of 64 for our evaluation.
@@
w/ Zero-shot Vocoder Choi et al. (2024) 23.76 0.154 3.088
w/ CosyVoice VC Du et al. (2024) 23.09 0.315 3.787
@@
Unit-to-speech module. To evaluate our unit-to-speech synthesizer’s ability to preserve speaker identity, we compare the performance with a zero-shot vocoder proposed in AV2AV Choi et al. (2024). Additionally, we applied a cross-lingual voice conversion model Du et al. (2024) to our single-speaker model and compare the results. As shown in Table 8, our unit-to-speech model achieves the highest translation quality, even outperforming the single-speaker model. While the zero-shot vocoder maintains reasonable translation quality, it shows poor speaker similarity. On the other hand, using a separate voice conversion model significantly degrades translation quality. On the other hand, using a separate voice conversion model shows a significant drop in ASR-BLEU. Based on qualitative analysis, we hypothesize that the process of voice conversion introduces oversmoothing of pronunciation that leads to loss in intelligibility and ultimately translation quality.
@@ 6 Conclusion
In this paper, we introduce a S2ST system applicable for dubbing applications. The proposed model generates translated speech with accurate content while preserving the duration and speaking speed of the source. This is achieved by the unique design that can generate arbitrary duration, and speed adaptation that removes the discrepancy between source and target speech. Extensive experiments with systematic ablations demonstrate that Dub-S2ST outperforms the existing baselines and verify its applicability to cross-lingual dubbing.
@@ 7 Limitations
While the proposed model offers an effective approach to cross-lingual dubbing, it is trained on speech recorded in controlled environments. This suggests that training on larger and more diverse datasets may be required for application to in-the-wild scenarios. Moreover, the proposed model is trained with a sentence-level speech corpus, and therefore a segmentation process is needed before being utilized in other applications.

### v2

@@ 4.1 Dataset
CVSS-C jia2022cvss is a widely-used dataset for speech-to-speech translation, consisting of 21 languages to English translations where English speech is generated by a single-speaker TTS model. The proposed method is trained and evaluated with French-English (fr-en) subset, due to the abundance of samples compared to other language pairs. The fr-en subset contains 207,364 (train) / 14,759 (dev) / 14,759 (test) pairs of source and target speech samples, totaling 264 hours. The experiment utilizes train and dev split for training and validation, and test split for final evaluation.
@@ 4.2 Implementation Details
Preprocessing. Audio samples are resampled to 16kHz and preprocessed with Voice Activity Detection (VAD) tool [^1] https://github.com/snakers4/silero-vad to remove unnecessary silence and paralinguistic information at the beginning and end. The samples are then processed with pretrained mHuBERT lee2021textless and k-means clustering model to obtain discrete units [^2] https://github.com/facebookresearch/fairseq/blob/main/examples/speech_to_speech/docs/textless_s2st_real_data.md.
Architecture. To maintain consistency with prior works on speech-to-unit translation models, we design our model with 12 Conformer encoder layers and 6 DiT layers, totaling 61M parameters. The unit-to-speech synthesizer consists of 4 Down, Mid, and UpBlock, where each is a cascade of 1D Convolution and Transformer layer.
Training. The speech-to-unit translation model is trained with a total batch size of 3,200 seconds for 100k updates. We use 8 RTX A5000 GPUs for training, and the total training takes approximately 10 hours. We optimize the model with the AdamW loshchilov2019decoupled optimizer and applied a dropout rate of 0.3. The learning rate is warmed up for the first 10k steps to a peak of $1\times 10^{-3}$, and then decayed using an inverse square root schedule. We implement our approach using Fairseq ott2019fairseq. The unit-to-speech module is initialized from CosyVoice-300M du2024cosyvoice and fine-tuned with fixed learning rate of $1\times 10^{-4}$ for 200k steps using LRS3 afouras2018lrs3 dataset, an English multi-speaker corpus.
@@ 4.3 Evaluation Metrics
ASR-BLEU lee2022direct is a widely adopted metric for assessing S2ST quality. It measures translation quality by transcribing the generated speech using a pretrained ASR model baevski2020wav2vec and comparing it with the ground-truth text to compute BLEU post2018call  [^3] https://github.com/facebookresearch/fairseq/tree/ust/examples/speech_to_speech/asr_bleu.
@@
S2UT lee2022direct 24.54 3.784 0.036 3.922
CTC-S2UT fang2024ctc 24.51 3.785 0.037 3.908
UTUT kim2024textless 26.49 3.840 0.036 3.927
w/ Zero-shot Vocoder choi2024av2av 26.33 3.882 0.145 3.101
TranSpeech huang2023transpeech 18.03 - - -
DiffNorm tan2024diffnorm 19.53 - - -
@@
BLASER 2.0 dale2024blaser serves as an automatic measure for assessing semantic similarity between source and generated speech. We adopt its reference-free variant, BLASER 2.0-QE [^4] https://huggingface.co/facebook/blaser-2.0-qe, which estimates translation quality without reference text unlike BLEU score.
SIM evaluates the speaker similarity between the generated and original speech. We use a pretrained speaker verification model chen2022large, based on WavLM-Large chen2022wavlm, to extract speaker embedding vectors and calculate cosine similarity between the two.
DNSMOS (Deep Noise Suppression Mean Opinion Score) reddy2021dnsmos is an automated perceptual speech quality assessment of generated speech [^5] We use a model trained with ITU-T P.808: https://github.com/microsoft/DNS-Challenge/tree/master/DNSMOS. The quality is estimated with a score in the range of [1, 5] where the larger value indicates higher speech quality.
Duration Compliance (DC) wu2023videodubber calculates the portion of generated speech whose duration ratio with the source lies within certain range. This indicates how S2ST system preserves the duration when generating translated speech.
Speed Compliance (SC) captures a similar aspect to DC but is based on the ratio between the speed of speech, measured in syllables per second [^6] https://github.com/facebookresearch/seamless_communication/blob/main/docs/expressive. This metric reflects how closely the speed of the generated speech aligns with that of the source.
@@ 5.1 Quantitative Comparison
Duration and Speed Analysis. We evaluate our model’s performance in duration control by measuring the generated speech duration relative to the source, as shown in Table 2. In addition to standard S2ST baselines, we compare against methods that incorporate duration control: VideoDubber wu2023videodubber, which employs additional positional embeddings, and TransVIP le2024transvip, which introduces isochrony positional embedding based on voice activity information. While existing methods achieve reasonable compliance within a 40% threshold, they struggle under a stricter 20% constraint. In contrast, our model achieves exact duration matching through explicit length initialization in the decoder, yielding 100% compliance.
The proposed model’s advantage becomes more evident in the speed analysis, where our model, both with and without speed adaptation, outperforms all baselines in speed compliance. This indicates that, when conditioned on the source duration, our model not only matches the duration but also the speaking pace of the source. The proposed speed adaptation further enhances this alignment. Overall, these results underscore the robustness of our approach in controlling speech duration and speed, both of which serve as critical factors in dubbing.
@@ 5.2 Human Evaluation
The speed consistency evaluation confirms that our proposed speed adaptation method effectively aligns the speaking speed between source and target speech, producing translated speech at a pace closely matching the source. This implicit synchronization contributes to improving the translation consistency, since it helps the model generate sequences at a similar speed to the source when given its duration, minimizing unnecessary pauses and repetitive words. Notably, significant differences in speech naturalness emerged during the evaluation. Baseline samples required manual waveform adjustments, leading to considerable degradation in perceptual quality. Conversely, our proposed approach yielded accurate, time-aligned speech outputs, eliminating the need for post-processing adjustments and maintaining high naturalness suitable for dubbing applications.
@@ 5.3 Discussion
Can explicit duration control truly change the translated content? As illustrated in Figure 3, our model responds to different duration prompts by producing semantically consistent outputs with nuanced variations in structure and phrasing. For example, “none of this happened” is rendered as “nothing took place” at 80% duration and “nothing of all that took place” at 120%, each expressing the same meaning without repetition or loss of information. Moreover, the relative number of characters of the translated outputs with different durations in Table 5 validates that the model flexibly generates semantics that fit the given duration, rather than forcing a fixed translation to fit varying lengths.
@@ 5.4 Ablation Study
Model architecture and Masking schedule. To assess how different model architectures and masking schedules affect the speech-to-unit translation model of Dub-S2ST, we compare several design choices, as shown in Table 6. The last row presents ASR-BLEU performances from Dub-S2ST-single without speed adaptation, evaluated across varying numbers of function evaluations (NFE). Replacing our DiT decoder with a standard Transformer decoder consistently degrades translation quality across all NFEs, indicating that incorporating diffusion timestep into the model benefits translation learning. We also examine the effect of the masking schedule by comparing our linear schedule with a cosine schedule. The results show that the linear schedule yields overall better performance, suggesting its effectiveness without losing translation quality. Based on the latency-performance trade-off, we choose the NFE of 64 for our evaluation.
@@
w/ Zero-shot Vocoder choi2024av2av 23.76 0.154 3.088
w/ CosyVoice VC du2024cosyvoice 23.09 0.315 3.787
@@
Unit-to-speech module. To evaluate our unit-to-speech synthesizer’s ability to preserve speaker identity, we compare the performance with a zero-shot vocoder proposed in AV2AV choi2024av2av. Additionally, we applied a cross-lingual voice conversion model du2024cosyvoice to our single-speaker model and compare the results. As shown in Table 8, our unit-to-speech model achieves the highest translation quality, even outperforming the single-speaker model. While the zero-shot vocoder maintains reasonable translation quality, it shows poor speaker similarity. On the other hand, using a separate voice conversion model shows a significant drop in ASR-BLEU. Based on qualitative analysis, we hypothesize that the process of voice conversion introduces oversmoothing of pronunciation that leads to loss in intelligibility and ultimately translation quality.
@@ 6 Conclusion
In this paper, we introduce an S2ST system suitable for dubbing applications. The proposed model generates translated speech with accurate content while preserving the duration and speaking speed of the source. This is achieved by the unique design that can generate speech with arbitrary duration, and speed adaptation that mitigates discrepancies between source and target speech. Extensive experiments with systematic ablations demonstrate that Dub-S2ST outperforms the existing baselines and verify its applicability to cross-lingual dubbing.
@@ 7 Limitations
While the proposed model offers an effective approach to cross-lingual dubbing, it is trained on speech recorded in controlled environments. This suggests that training on larger and more diverse datasets may be required for application to in-the-wild scenarios. Moreover, the proposed model is trained on a sentence-level speech corpus, and therefore a segmentation process is required before being utilized in other applications.
Acknowledgments
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grants funded by the Korean government (MSIT, RS-2025-02215122, Development and Demonstration of Lightweight AI Model for Smart Homes and RS-2022-II220989, Development of Artificial Intelligence Technology for Multi-speaker Dialog Modeling).